The college Physical Education Department offered an Advanced First Aid course last summer. The scores on the comprehensive final exam were normally distributed, and the z scores for some of the students are shown below.
Robert, 1.03 Juan, 1.89 Susan, –1.99
Joel, 0.00 Jan, –0.69 Linda, 1.46
(d) If the mean score was μ = 159 with standard deviation σ = 25, what was the final exam score for each student? (Round your answers to the nearest whole number.) Robert
Joel
Jan
Juan
Susan
Linda

Answer by stanbon(75887)

The college Physical Education Department offered an Advanced First Aid course last summer. The scores on the comprehensive final exam were normally distributed, and the z scores for some of the students are shown below.
Robert, 1.03 Juan, 1.89 Susan, –1.99
Joel, 0.00 Jan, –0.69 Linda, 1.46
(d) If the mean score was μ = 159 with standard deviation σ = 25, what was the final exam score for each student? (Round your answers to the nearest whole number.)
-----
Use x = z*s + u
-------------------
Robert
x = 1.03*25 + 159 = 184.75
-----
Use the same procedure for the other names.
=============
Cheers,
Stan H.
=============